OpenStudy (frostbite):

About the steady state approximation ((BIO)chemical kinetics) @shrutipande9

OpenStudy (shrutipande9):

ohh yes please....see i just need to clear d concept....i dnt want to go into d "chemistry" details...

OpenStudy (frostbite):

Ok I am going to summarize the principle in one sentence: "The steady-state principle assumes a constant but dynamic concentration of reaction intermediates"

OpenStudy (frostbite):

Ok what does mean: Consider a reaction \[\huge \sf A \overset{k_{1}}{\rightarrow} B \overset{k_{2}}{\rightarrow} C\] From a avg reaction we could imagine the cures to look like this:  Ok so what we see is that the intermediate is being created and then faints away as C is being produced. Lets now assume that \(\Large \sf k_{2}>k_{1} \) we then get to see for a enzymatic reaction:  We see that B remains constant after a little time (we call that pre-steady-state). Mathematical we get to say: \[\Large \frac{ d[B] }{ dt }=k_{1} \times [A]-k_{2} \times [B]=0\] Where the differential is an expression for the rate. By this approximation we get to see that: \[\Large k_{1} \times [A]=k_{2} \times [B]\] This expression is very useful to evaluate the rate equations!